### v1

@@ 6 Conclusion
In this paper, we derive the sharp testable implications for identifying assumptions for the judge’s leniency design in a general framework where the instruments can be either discrete or continuous and propose a consistent test for the implications. Our simulation study and empirical results identify how essential it is to consider sharp implications for the best use of information in the data. While we focus on the primary application of testing the validity of judge leniency design, our method can readily be applied to a broad range of applications.
@@ Appendix A Implementation of the test
In this section, we describe the details of calculating the test statistics for 3.1. Let $\{Y_{i},Z_{i},D_{i}\}_{i=1}^{n}$ be a random sample and $P(D_{i}=1|Z_{i})=P(Z_{i},\theta_{0})$ be the propensity known to $\theta_{0}$.
@@ A.1 Constructing $\hat{\nu}_{d}(\ell)$
First, note that
@@
$\hat{m}_{d}(y,r_{y},p,r_{p},\hat{\theta})=\frac{1}{n}\sum_{i=1}^{n}m_{di}(y,r_{y},p,r_{p},\hat{\theta}),\quad d=0,1$ (A.4)
$\hat{w}(p,r_{p},\hat{\theta})=\frac{1}{n}\sum_{i=1}^{n}w_{i}(p,r_{p},\hat{\theta}).$ (A.5)
@@
$\hat{\nu}_{1}(\ell)=\hat{m}_{1}(y,r_{y},p_{2},r_{p},\hat{\theta})\cdot\hat{w}(p_{1},r_{p},\hat{\theta})-\hat{m}_{1}(y,r_{y},p_{1},r_{p},\hat{\theta})\cdot\hat{w}(p_{2},r_{p},\hat{\theta}),$ (A.6)
$\hat{\nu}_{0}(\ell)=\hat{m}_{0}(y,r_{y},p_{2},r_{p},\hat{\theta})\cdot\hat{w}(p_{1},r_{p},\hat{\theta})-\hat{m}_{0}(y,r_{y},p_{1},r_{p},\hat{\theta})\cdot\hat{w}(p_{2},r_{p},\hat{\theta}).$ (A.7)
@@ A.2 Constructing $\hat{\nu}_{d}^{b}(\ell)$
For bootstrap iteration $b$, let $\{W_{1}^{b},W_{2}^{b},\cdots,W_{n}^{b}\}$ be a sequence of i.i.d. random variables with zero mean and unit variance. For instance, we can choose standard normal. Let $\hat{\theta}^{b}$ be the MLE based on the $b$-th bootstrapped sample. We define the weighted bootstrapped estimators for $m_{1}(y,r_{y},p,r_{p},\theta_{0})$, $m_{0}(y,r_{y},p,r_{p},\theta_{0})$ and $w(p,r_{p},\theta_{0})$ be
@@
A Sharp Test for the Judge Leniency Design
Online Supplementary Materials
Mohamed Coulibaly     Yu-Chin Hsu    Ismael Mourifié     Yuanyuan Wan
Appendix B Proof of Theorem 2.1
@@ A Sharp Test for the Judge Leniency Design
$\Omega(\mathcal{P})\equiv\big{\{}(p,p^{\prime}]:p,p^{\prime}\in I(\mathcal{P})\cup\{0,1\}\text{ and for all $\tilde{p}$ such that }p<\tilde{p}<p^{\prime},\tilde{p}\notin\mathcal{P}\big{\}}.$
So the interior of each interval does not intersect with $\mathcal{P}$. $\Omega(\mathcal{P})$ contains a generic element $(c_{k},d_{k}]$, where $c_{k}$, $d_{k}\in I(\mathcal{P})$, $d_{k}\leq c_{k+1}$, $k=1,2,\cdots,K$ with $K$ possibly equals to $\infty$, depending on how many isolation points there are in $\mathcal{P}$. Note that with above notation, for any $v\in(0,1]$, $v$ must belongs to one of the following categories: (i) an element of $\mathcal{L}^{o}(\mathcal{P})$ so that $v\in(a_{j},b_{j})$ for some $j$, (ii) $v\in\mathcal{L}(\mathcal{P})/\mathcal{L}^{o}(\mathcal{P})$, and (iii) there exist an integer $k$ such that $v\in(c_{k},d_{k}]$. The following figure illustrates the partition of the unit interval.
@@
$\mathbb{P}(\tilde{V}\leq t|P=p)=t,\forall(t,p)\in[0,1]\times\mathcal{P},\text{and }\tilde{D}=1\{P(Z)\geq\tilde{V}\}.\$
@@
$\mathbb{P}(\tilde{Y}_{1}\leq y|\tilde{V}=v,P=p)$ $=$ $\begin{cases}\frac{\partial}{\partial t}\mathbb{P}(Y\leq y,D=1|P=t)|_{t=v}\;\;\text{ if }v\in\mathcal{L}^{o}(\mathcal{P})\\ \lim_{\tilde{v}\rightarrow v}\frac{\partial}{\partial t}\mathbb{P}(Y\leq y,D=1|P=t)|_{t=\tilde{v}}\;\;\text{ if }v\in\mathcal{L}(\mathcal{P})/\mathcal{L}^{o}(\mathcal{P})\\ \frac{\mathbb{P}(Y\leq y,D=1|P=d_{k})-\mathbb{P}(Y\leq y,D=1|P=c_{k})}{d_{k}-c_{k}}\;\;\text{ if }v\notin L(P)\text{ but }v\in(c_{k},d_{k}]\in\Omega(\mathcal{P}).\\ \end{cases}$
$\mathbb{P}(\tilde{Y}_{0}\leq y|\tilde{V}=v,P=p)$ $=$ $\begin{cases}-\frac{\partial}{\partial v}\mathbb{P}(Y\leq y,D=0|P=t)|_{t=v}\;\;\text{ if }v\in\mathcal{L}^{o}(\mathcal{P})\\ -\lim_{\tilde{v}\rightarrow v}\frac{\partial}{\partial v}\mathbb{P}(Y\leq y,D=0|P=t)|_{t=\tilde{v}}\;\;\text{ if }v\in\mathcal{L}^{o}(\mathcal{P})\\ \frac{\mathbb{P}(Y\leq y,D=0|P=c_{k})-\mathbb{P}(Y\leq y,D=0|P=d_{k})}{d_{k}-c_{k}}\;\;\text{ if }v\notin L^{o}(P)\text{ but }v\in(c_{k},d_{k}]\in\Omega(\mathcal{P}).\\ \end{cases}$
@@
    $\mathbb{P}(\tilde{Y}_{1}\leq y^{\prime}|\tilde{V}=v,P=p)-\mathbb{P}(\tilde{Y}_{1}\leq y|\tilde{V}=v,P=p)\\ =\begin{cases}\frac{\partial}{\partial t}\mathbb{P}(y<Y\leq y^{\prime},D=1|P=t)|_{t=v}\geq 0\;\;\text{ if }v\in\mathcal{L}^{o}(\mathcal{P}),\\ \lim_{\tilde{v}\rightarrow v}\frac{\partial}{\partial t}\mathbb{P}(y<Y\leq y,D=1|P=t)|_{t=\tilde{v}}\geq 0\;\;\text{ if }v\in\mathcal{L}(\mathcal{P})/\mathcal{L}^{o}(\mathcal{P})\\ \frac{\mathbb{P}(y<Y\leq y^{\prime},D=1|P=d_{k})-\mathbb{P}(y<Y\leq y^{\prime},D=1|P=c_{k})}{d_{k}-c_{k}}\geq 0\text{ if }v\notin L^{o}(P)\text{ but }v\in[c_{k},d_{k}]\in\Omega(\mathcal{P}),\end{cases}$
@@
Finally, we show that $(\tilde{V},\tilde{Y}_{d},P(Z))$, $d\in\{0,1\}$ is observationally equivalent to $(V,Y_{d},P(Z))$ $d\in\{0,1\}$. For this, we show that the conditioning distribution of $(\tilde{Y},\tilde{D})$ given $P(Z)$ is the same as the conditioning of $(Y,D)$ given $P(Z)$. Take an arbitrary $p\in\mathcal{P}$.
@@
where the $\int_{0}^{a_{j^{*}}}\mathbb{P}(\tilde{Y}_{1}\leq y|\tilde{V}=v,P=p)dv=\mathbb{P}(Y\leq y,D=1|P=a_{j^{*}})$ holds by the above argument and the fifth equality holds by inserting the constructed counterfactural distributions. This completes the $\boldsymbol{\Box}$
Appendix C Proof of Theorem 3.2
We begin by listing a few regularity conditions for the proof of Theorem 3.2.
Assumption C.1
@@
Assumption C.2
@@
Note that Assumption C.2-(1) does not exclude the case of discrete propensity score. When $P$ is discrete and $\mathcal{P}$ contains finite many distinguished elements, any convergent sequence in $\mathcal{P}$ must be a constant sequence eventually, and in that case Assumption C.2-(1) holds automatically. Assumption C.2-(1) implies that the functions $m_{d}$ and $\omega$, defined in Equations A.1, A.2 and A.3, are continuous functions of $\ell$. Assumption C.2-(2) implies that the class of functions $\{1(p\leq P(Z,\theta)\leq p+r_{p}):\theta\in\Theta,p\in[0,1],r_{p}\in[0,1]\}$ is a Vapnik-Chervonenkis (VC) class of function.
Assumption C.3
@@
$\sqrt{n}(\hat{\theta}-\theta_{0})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}s(D_{i},Z_{i},\theta_{0})+o_{p}(1),$ (C.1)
@@
Assumption C.3 is satisfied for common maximum likelihood estimators and parametric binary response models. For example, if one estimates $\theta_{0}$ by Probit model $D_{i}=1[Z_{i}^{\prime}\theta_{0}\geq V_{i}]$, with $V_{i}\sim N(0,1)$, then
@@
Assumption C.4
@@
$\sqrt{n}(\hat{\theta}^{b}-\hat{\theta})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(W_{i}-1)\cdot s(D_{i},Z_{i},\theta_{0})+o_{p}(1),$ (C.2)
where $s_{\theta}(\cdot)$ is the same as in Assumption C.3.
Assumption C.4 is satisfied under our weighted bootstrap procedure.

### v2

@@ 6 Conclusion
In this paper, we derive the sharp testable implications for identifying assumptions for the judge’s leniency design in a general framework where the instruments can be either discrete or continuous and propose a consistent test for the implications. Our simulation study and empirical results highlight the importance of considering sharp implications for a better use of information in the data. While we focus on the primary application of testing the validity of judge leniency design, our method can be readily applied to a broad range of other applications.
@@ Appendix A Implementation of the test
In this section, we describe the details of calculating the test statistics for 3.1. Let $\{Y_{i},Z_{i},D_{i}\}_{i=1}^{n}$ be a random sample and $P(D_{i}=1|Z_{i})=P(Z_{i},\theta_{0})$ be the propensity known to $\theta_{0}$. Note that when $Z$ is the judge’s identity and if the number of defendants for each judge diverges to infinity, we can simply use the frequency estimator $\hat{P}_{i}=\frac{\sum_{k=1}^{n}D_{k}1\{Z_{k}=Z_{i}\}}{\sum_{k=1}^{n}1\{Z_{k}=Z_{i}\}}$. Therefore, in the appendix sections, we focus on the case in which $Z$ is continuous to simplify notation.
@@ A.1 Constructing $\hat{\nu}_{d}(\ell)$
First, when $Z$ is continuous, we estimate $\theta_{0}$ by MLE,
$\equiv\mathop{\mathrm{argmax}}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}D_{i}\log P(Z_{i},\theta)+(1-D_{i})\log(1-P(Z_{i},\theta)).$ (A.1)
where $P(z,\theta)$ is parameterized and depends on $z$ through $z^{\prime}\theta$. For example, $P(z,\theta)=\Phi(z^{\prime}\theta_{z})$ for Probit or $P(z,\theta)=\frac{exp(z^{\prime}\theta_{z})}{1+\exp(z^{\prime}\theta_{z})}$ for Logit.
Next, note that
@@
$\hat{m}_{d}(y,r_{y},p,r_{p},\hat{\theta})=\frac{1}{n}\sum_{i=1}^{n}m_{di}(y,r_{y},p,r_{p},\hat{\theta}),\qquad d=0,1$ (A.5)
[…]
@@
[…]
@@ A.2 Constructing $\hat{\nu}_{d}^{b}(\ell)$
In this appendix, we show how to construct the bootstrap estimates $\hat{\nu}_{d}^{b}(\ell)$. For bootstrap iteration $b$, let $\{W_{1}^{b},W_{2}^{b},\cdots,W_{n}^{b}\}$ be a sequence of i.i.d. random variables with both mean and variance equal to one. For instance, we can choose standard normal. Let $\hat{\theta}^{b}$ be the MLE based on the $b$-th bootstrapped sample:
$\hat{\theta}^{b}=\mathop{\mathrm{argmax}}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}W_{i}^{b}\log f(Y_{i},D_{i};Z_{i}\theta)\\ \equiv\mathop{\mathrm{argmax}}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}W_{i}^{b}\left\{D_{i}\log P(Z_{i},\theta)+(1-D_{i})\log(1-P(Z_{i},\theta))\right\},$
and the estimated propensity score for the $b$-th bootstrap as
$\hat{P}_{i}^{b}=P(Z_{i},\hat{\theta}^{b})$ (A.9)
We define the weighted bootstrapped estimators for $m_{1}(y,r_{y},p,r_{p},\theta_{0})$, $m_{0}(y,r_{y},p,r_{p},\theta_{0})$ and $w(p,r_{p},\theta_{0})$ be
@@
Appendix B Proof of Main Results
B.1 Proof of Theorem 2.1
@@ Appendix B Proof of Main Results
$\Omega(\mathcal{P})\equiv\big\{(p,p^{\prime}]:p,p^{\prime}\in I(\mathcal{P})\cup\{0,1\}\text{ and }(p,p^{\prime})\cap\mathcal{P}=\emptyset\big\}.$
So the interior of each interval does not intersect with $\mathcal{P}$. $\Omega(\mathcal{P})$ contains a generic element $(c_{k},d_{k}]$, where $c_{k}$, $d_{k}\in I(\mathcal{P})$, $d_{k}\leq c_{k+1}$, $k=1,2,\cdots,K$ with $K$ possibly equals to $\infty$, depending on how many isolation points there are in $\mathcal{P}$. Note that with above notation, for any $\in(0,1], vv$ must belongs to one of the following categories: (i) an element of $\mathcal{L}^{o}(\mathcal{P})$ so that $v\in(a_{j},b_{j})$ for some $j$, (ii) $v\in\mathcal{L}(\mathcal{P})/\mathcal{L}^{o}(\mathcal{P})$, and (iii) there exist an integer $k$ such that $v\in(c_{k},d_{k}]$. The following figure illustrates the partition of the unit interval.
@@
$\mathbb{P}(\tilde{V}\leq t|P=p)=t,\forall(t,p)\in[0,1]\times\mathcal{P},\text{ and }\tilde{D}=1\{P(Z)\geq\tilde{V}\}$.\\
@@
[…]
@@
[…]
@@
[…]
@@
$\mathbb{P}(\tilde{Y}\leq y,\tilde{D}=1|P=p)=\mathbb{P}(\tilde{Y}_{1}\leq y,\tilde{V}\leq p|P=p)=\int_{0}^{p}\mathbb{P}(\tilde{Y}_{1}\leq y|\tilde{V}=v,P=p)dv\\ =\int_{0}^{a_{j^{*}}}\mathbb{P}(\tilde{Y}_{1}\leq y|\tilde{V}=v,P=p)dv+\int_{a_{j^{*}}}^{p}\mathbb{P}(\tilde{Y}_{1}\leq y|\tilde{V}=v,P=p)dv\\ =\mathbb{P}(Y\leq y,D=1|P=a_{j^{*}})+\int_{a_{j^{*}}}^{p}\frac{\partial}{\partial v}\mathbb{P}(Y\leq y,D=1|P=v)dv\\ =\mathbb{P}(Y\leq y,D=1|P=a_{j^{*}})+\mathbb{P}(Y\leq y,D=1|P=p)-\mathbb{P}(Y\leq y,D=1|P=a_{j^{*}})\\ =\mathbb{P}(Y\leq y,D=1|P=p),$
@@
B.2 Proof of Theorem 3.2
We begin by listing a few regularity conditions for the proof of Theorem 3.2. Again, when $Z$ is the judge’s identity, we use the frequency estimator $\hat{P}_{i}=\frac{\sum_{k=1}^{n}D_{k}1\{Z_{k}=Z_{i}\}}{\sum_{k=1}^{n}1\{Z_{k}=Z_{i}\}}$ for the propensity score. For its root-n-consistency, we only need $\sum_{i=1}^{n}1\{Z_{k}=j\}\rightarrow\infty$ for each judge $j$, and i.i.d. of $(Y_{i},D_{i},X_{i})$ among defendants conditioning on judges. So Assumptions B.1, B.2, B.3 and B.5 are mostly for the case of continuous instrument $Z$.
Assumption B.1
@@
For notational simplicity, Assumption B.1 assumes that all cases are mutually independent, which is equivalent to assuming that each judge handles exactly one case. All inference results can be extended straightforwardly to settings where judges handle multiple cases (with varying case counts across judges) by accounting for clustering at the judge level.
Assumption B.2
@@
Note that Assumption B.2-(1) does not exclude the case of discrete propensity score. When $P$ is discrete and $\mathcal{P}$ contains finite many distinguished elements, any convergent sequence in $\mathcal{P}$ must be a constant sequence eventually, and in that case Assumption B.2-(1) holds automatically. Assumption B.2-(1) implies that the functions $m_{d}$ and $\omega$, defined in Equations A.2, A.3 and A.4, are continuous functions of $\ell$. Assumption B.2-(2) implies that the class of functions $\{1(p\leq P(Z,\theta)\leq p+r_{p}):\theta\in\Theta,p\in[0,1],r_{p}\in[0,1]\}$ is a Vapnik-Chervonenkis (VC) class of function.
Assumption B.3
@@
$\sqrt{n}(\hat{\theta}-\theta_{0})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}s(D_{i},Z_{i},\theta_{0})+o_{p}(1),$ (B.1)
@@
Assumption B.3 is satisfied for common maximum likelihood estimators and parametric binary response models. For example, if one estimates $\theta_{0}$ by Probit model $D_{i}=1[Z_{i}^{\prime}\theta_{0}\geq V_{i}]$, with $V_{i}\sim N(0,1)$, then
@@
Assumption B.4
$\{W_{i}\}_{i=1}^{n}$ is a sequence of i.i.d. pseudo random variables that is independent of the sample path with $E[W_{i}]=1$ and $Var[W_{i}]=1$.
Assumption B.5
@@
$\sqrt{n}(\hat{\theta}^{b}-\hat{\theta})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(W_{i}-1)\cdot s(D_{i},Z_{i},\theta_{0})+o_{p}(1),$ (B.2)
where $s_{\theta}(\cdot)$ is the same as in Assumption B.3.
Assumption B.5 is satisfied under our weighted bootstrap procedure.